OpenStudy (anonymous):

Find formula for upper sum, then take limit of sums to calculate area: f(x)=3x^2 +x over the interval [0,1] sum_{i=1}^{n} i= n(n+1)/2, sum_{i=1}^{n}= n(n+1)(2n+1)/6

OpenStudy (zarkon):

notice that on [0,1] f(x) is increasing so the max on any sub interval is the right hand endpoint

OpenStudy (zarkon):

I'll try my best...thanks for the reminder :)

OpenStudy (zarkon):

lol

OpenStudy (anonymous):

ok, so i know the correct format would be:\[\sum_{i=1}^{n} i= n(n+1)/2, \sum_{i=1}^{n}= n(n+1)(2n+1)/6\sum\]

OpenStudy (anonymous):

uh oh, ignore that last sigma

OpenStudy (zarkon):

you forgot your \(i^2\) again

OpenStudy (anonymous):

oh no, ok i'll try again :(

OpenStudy (anonymous):

\[\sum_{i=1}^{n} i= n(n+1)/2, \sum_{i=1}^{n}i^2= n(n+1)(2n+1)/6\]

OpenStudy (zarkon):

better :)

OpenStudy (anonymous):

so what steps would I have to actually do to solve the rest of the problem?

OpenStudy (zarkon):

you need to calculate \[\sum_{i=1}^{n}f(x_i^*)\Delta x\]

OpenStudy (zarkon):

in this case you can use \(x_i^*=\frac{i}{n}\) and \(\Delta x=\frac{1}{n}\)

OpenStudy (anonymous):

would it be too much to ask if you showed me how to do this problem step by step? I learn best by example :)

OpenStudy (zarkon):

do you understand what I have done so far?

OpenStudy (anonymous):

i think so

OpenStudy (zarkon):

we then need to compute \[\sum_{i=1}^{n}f(x_i^*)\Delta x=\sum_{i=1}^{n}f\left(i\frac{1}{n}\right)\frac{1}{n}\]

OpenStudy (zarkon):

\[=\sum_{i=1}^{n}\left(3\left(i\frac{1}{n}\right)^2+i\frac{1}{n}\right)\frac{1}{n}\]

OpenStudy (zarkon):

ok?

OpenStudy (anonymous):

yep, makes sense so far

OpenStudy (zarkon):

\[=\sum_{i=1}^{n}\left(3\left(i^2\frac{1}{n^2}\right)+i\frac{1}{n}\right)\frac{1}{n}\] \[=\sum_{i=1}^{n}\left(3i^2\frac{1}{n^3}+i\frac{1}{n^2}\right)\]

OpenStudy (zarkon):

\[=3\frac{1}{n^3}\sum_{i=1}^{n}i^2+\frac{1}{n^2}\sum_{i=1}^{n}i\]

OpenStudy (zarkon):

\[=3\frac{1}{n^3}\frac{n(n+1)(2n+1)}{6}+\frac{1}{n^2}\frac{n(n+1)}{2}\]

OpenStudy (anonymous):

oh, is that the final answer?

OpenStudy (zarkon):

no

OpenStudy (anonymous):

its just the formula to solve for the area right?

OpenStudy (zarkon):

\[\lim_{n\to\infty}\left[3\frac{1}{n^3}\frac{n(n+1)(2n+1)}{6}+\frac{1}{n^2}\frac{n(n+1)}{2}\right]=1+\frac{1}{2}=\frac{3}{2}\]

OpenStudy (anonymous):

so 3/2 would be the answer?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thank you. would you mind helping me with an integral problem?

OpenStudy (zarkon):

I have to go...Will probably be back on later. Post your problem in a new thread and I'm sure someone will help

OpenStudy (anonymous):

ok, well thanks anyways :)